### v1

# Emergent photons and mechanisms of confinement

Abstract
We numerically study $\mathbb{Z}_{N}$ lattice gauge theories in 4D as prototypical models of systems with $\mathbb{Z}_{N}$ 1-form symmetry. For $N\geq 3$, we provide evidence that such systems exhibit not only the expected phases with spontaneously broken/restored symmetry but also a third photon phase. When present, the 1-form symmetry provides a precise notion of confinement, and it is commonly believed that confinement ensues due to the proliferation of extended, string-like objects known as center vortices, which carry a $\mathbb{Z}_{N}$ flux. However, this picture is challenged by the three-phase scenario investigated here. We show that both the confined and the photon phases are associated with the proliferation of center vortices and that the key difference between them lies in whether or not vortex-junctions — the monopoles — proliferate.
I Introduction
At center stage of the Standard Model of particle physics, two theories play a distinguished role: the theory of Quantum Electrodynamics (QED), describing light and chemistry, and the theory of Quantum Chromodynamics (QCD), describing the formation of nuclei. While the former is a theory of massless photons that enables us to see far into our universe, the latter is a system with a tiny correlation length of $\sim 1$ fm and is responsible for most of the atomic mass.
Yet both of these theories and regimes are linked to an exotic symmetry called the 1-form symmetry that has been of particular interest in the recent decade, mostly starting with the works [1, 2, 3] [^1] The 1-form symmetries were known in $\text{SU}(N)$ Yang-Mills theory at least since the late 70s [4], but until recently were relatively poorly understood.. On one hand, in QCD, quark confinement is associated with the would-be $Z_{N}$ 1-form symmetry — a symmetry of the pure glue sector — while the massless photons of QED are seen as Goldstone bosons of a $U(1)$ 1-form symmetry. [2, 5]. Therefore, both the massless photon regime and the nuclear confined regime can potentially be understood as phases of matter within the Ginzburg-Landau paradigm for 1-form symmetries (see also [6]).
In [7], it was argued that theories with $\mathbb{Z}_{N}$ 1-form symmetry in 3+1D have a generic photon phase for any $N$. This phase is seen as an enhancement of the 1-form symmetry $\mathbb{Z}_{N}\rightarrow\text{U}(1)$, and consequent spontaneous breaking of it, leading to a massless Goldstone boson — the photon.
In this work, we will consider $\mathbb{Z}_{N}$ lattice gauge theory (LGT), a model of classical Ising-like spins on the links of the lattice. These theories have been known to exhibit three phases when $N\geq 5$ [8, 9]. Here, we show numerically that this is also true for $N=3,4$.
In parallel, the 3+1D $\mathbb{Z}_{N}$ LGT is a toy model of pure $SU(N)$ Yang-Mills theory, having both confined and deconfined phases. This model offers a deceptively simple picture of confinement as originating from the proliferation of string-like objects analogous to center vortices of $\text{SU}(N)$ Yang-Mills theory (see [10, 11] for a review). However, the existence of the photon phase requires this picture to be revised by also considering monopole particles [12, 7].
It is worth noting that the occurrence of a robust massless phase without a continuous global symmetry is surprising from the perspective of ordinary 0-form symmetries, as there can be no Goldstone bosons [^2] Alternatively, one could hit a conformal fixed point, but this typically requires fine-tuning.. Moreover, continuous symmetry cannot simply emerge, as any explicit breaking of an emergent 0-form continuous symmetry will lift the mass of the Goldstone boson.
That the same conclusion does not hold for $\mathbb{Z}_{N}$ 1-form symmetry in 3+1D is what makes the Millennium Problem of whether $\text{SU}(N)$ Yang-Mills theory has a mass gap [13] nontrivial. Indeed, based on the $\mathbb{Z}_{N}$ 1-form global symmetry group, we may be erroneously tempted to conclude that a gapless phase is untenable. Incidentally, in 2+1D $\text{SU}(N)$ Yang-Mills theory, the photon phase is generically unstable [7], resulting in the necessity of a mass gap under all generic symmetry-preserving deformations (see e.g. [14, 15, 16, 17]).
II The $\mathbb{Z}_{N}$ Lattice Gauge Theory (LGT)
Figure 1: a) The definition of the plaquette Wilson loop $W_{p}=U_{\ell_{1}}U_{\ell_{2}}U_{\ell_{3}}U_{\ell_{4}}$. b) The illustration of the $\mathbb{Z}_{N}$ 1-form symmetry transformation in 2D. c) The center vortex as an excitation of the plaquette fluxes.
The model we consider here is a $Z_{N}$ Lattice Gauge Theory, which is to $Z_{N}$ 1-form symmetry what the Ising model is to an ordinary $Z_{2}$ symmetry. Like the Ising model, it is a theory with discrete $N$-state spins $s_{\ell}=0,1,2,\cdots,N-1$ on lattice links $\ell$. The spins $s_{\ell}$ naturally form a $\mathbb{Z}_{N}$ phase $U_{\ell}=e^{\frac{2\pi is_{\ell}}{N}}$. An action that dictates the dynamics of these spins can be constructed by defining $W_{p}=\prod_{\ell\in\partial p}U_{\ell}$ — a plaquette sized Wilson loop — where $\partial p$ is the set of boundary links of a plaquette $p$, inheriting the orientation of the plaquette (with the convention $U_{\ell}=U_{-\ell}^{*}$). Explicitly, consider a hyper-cubic lattice, and let $\ell_{1},\ell_{2},\ell_{3},\ell_{4}$ be the boundary links of $p$ as in Fig. 1, then $W_{p}=U_{\ell_{1}}U_{\ell_{2}}U_{\ell_{3}}U_{\ell_{4}}$. The standard Wilson action and the partition function are given by
$S=-\beta\sum_{p}\text{Re\;}{W_{p}}\;,\quad Z=\sum_{\{s\}}e^{-S}\;,$ (1)
where the sum is over all spin configurations $\{s_{\ell}\}$.
A crucial property of this action in general dimension $d$ is that it is invariant under the $\mathbb{Z}_{N}$ 1-form symmetry [2]. To understand this symmetry applied to (1), consider an oriented $(d-1)$-dimensional hypersurface $H$ constructed out of $(d-1)$-cells of the dual lattice (an example for $d=2$ is given in Fig. 1). This will intersect a set of lattice links, which we collect into a set $L$ (marked blue in Fig. 1) and orient compatibly relative to $H$. A $\mathbb{Z}_{N}$ 1-form symmetry is generated by shifting [^3] We note that by convention $s_{-\ell}=-s_{\ell}\bmod N$. $s_{\ell}\rightarrow s_{\ell}+1\bmod N$ whenever $\ell\in L$; this is clearly an invariance of the action.
By taking $H$ to be such that it surrounds a single lattice site, the resulting transformation is the conventional gauge transformation.
Any contractible $H$ thus corresponds to a composition of local gauge transformations. On the other hand, non-contractible choices of $H$ that differ by a pure boundary yield the same transformation up to gauge symmetry. Such choices of $H$, up to pure boundary terms, are what is meant by a 1-form symmetry, so the global symmetry in question takes values in $H_{d-1}(\mathbb{Z}_{N})\cong H^{1}(\mathbb{Z}_{N})$.
The operators charged under the 1-form symmetry are Wilson loops $W[\mathcal{C}]=\prod_{\ell\in\mathcal{C}}U_{\ell}$ where $\mathcal{C}$ is a contour made out of the links of the lattice. It is easy to see that under the $1$-form symmetry transformation of Fig. 1b) the Wilson loop transforms as $W[\mathcal{C}]\rightarrow W[\mathcal{C}]\exp\left({i\frac{2\pi I(H,\mathcal{C})}{N}}\right)$ where $I(H,\mathcal{C})$ is the net intersection number between $H$ and $\mathcal{C}$. The fact that in the continuum the charged operator is a Wilson loop, which is an integral over the 1-form gauge field $A_{\mu}$ is what gives the 1-form symmetry its name.
The naive Landau paradigm would then indicate at least two phases: the $\mathbb{Z}_{N}$ restored (confined) and the spontaneously (symmetry) broken (SSB) phase, which can be seen respectively in the limits $\beta\rightarrow 0,\infty$. When $0<\beta\ll 1$, the contractible Wilson loops $W[\mathcal{C}]$ get a leading contribution at order $\beta^{A}$, where $A$ is the area of a minimal surface $S$ whose boundary is $\mathcal{C}$. By cluster decomposition, this implies that the non-contractible Wilson loops must have zero expectation value, i.e., the $\mathbb{Z}_{N}$ 1-form symmetry is restored. In the limit $\beta\gg 1$, the dominant configurations obey $W_{p}=1$, indicating that non-contractible Wilson loops get a nonzero expectation value.
Further, in this regime, excitations for which $W_{p}\neq 1$ (see Fig. 1c) can be thought of as heavy string-like excitations in 3+1D describing a 2D surface on the dual lattice. These string-like objects are analogous to center vortices studied in the QCD community, and this is the name we will adopt. One of the most important features of the center vortex is that whenever it links with a Wilson loop, it is weighted by a $\mathbb{Z}_{N}$ phase, giving a compelling picture of confinement that can be summarized as follows. As one lowers $\beta$, the center vortex tension goes down, and increasingly larger vortices become favored. At some value of $\beta$, the vortices proliferate, causing dramatic interference with the Wilson loop and eventual area law behavior. This is the essence of the so-called center vortex mechanism of confinement [18, 19].
It has been known since 1979 that for $N\geq 5$, the $\mathbb{Z}_{N}$ LGT in 4D has a robust phase of emergent photons for intermediate $\beta$ [8, 9], making it clear that center vortex proliferation alone cannot possibly explain both the photon and the confined phase. Therefore, a new ingredient is needed here [7]: the worldline representing a junction of $N$ center vortices — the monopole-junction.
Refer to caption Refer to captionRefer to captionRefer to caption
Figure 2: Phase structure of $\mathbb{Z}_{7}$ Lattice Gauge Theory. Plots show the average Polyakov loop and the corresponding distributions of the smeared Polyakov loop $\bar{P}$. The histograms of $\bar{P}$ indicate confinement (left), emergent U(1) symmetry (middle), and spontaneous $\mathbb{Z}_{7}$ 1-form symmetry breaking (right). Insets: Typical 3D Monte Carlo configurations showing center vortices (colored by charge ±1, ±2, ±3, arrows indicating the sign) and monopoles/anti-monopoles (green/yellow boxes).
In fact, a picture of the photon and the confined phase was first briefly touched on in Appendix B of the little-known paper by Michael Stone [12] and recently in [7], which we aim to confirm here. In this framework, both the photon and confined phases involve proliferating center vortices, but differ in the behavior of monopole-junctions: these remain heavy in the photon phase, while they proliferate in the confined phase. In this work, we confirm this idea.
We highlight two important observations:
  • The three phases known to exist for $N\geq 5$ are indeed associated with the (non-)proliferation of the center vortices and monopoles in the way described above and in [7]
  • The theories for $N=3,4$ have a deformation which allows for three phases
Conspicuously, we will not discuss the theory with $N=2$ in this work because it has no charge-conjugation symmetry and hence no natural choice of monopole-junction. The question of whether theories with (only) $\mathbb{Z}_{2}$ 1-form symmetry have a photon phase is left for future explorations.
III Numerical simulations and results
Refer to captionRefer to caption
Figure 3: Susceptibility $\chi_{S}$ versus $\beta$ for the $\mathbb{Z}_{3}$ model with monopole suppression ($\mu$=1) for lattices $L=8,10,12$ and $14$. Two prominent peaks indicate phase transitions. Bottom panels show zooms of the low-$\beta$ (confined $\to$ photon) and high-$\beta$ (photon $\to$ SSB) transition regions. Inset: smeared Polyakov loop histograms near the high-$\beta$ transition for $L=6,10$ and $14$.
In this section, we numerically study the two different classes of theories: $N\geq 5$ and $52<N<$ 5. We choose $N=7$ to represent the $N\geq 5$ regime in which three phases are well established [9, 8]. The goal is to associate these phases with (non-)proliferating center vortices and monopole-junctions. To this end, we begin by formally defining these objects on the lattice.
The center vortex is easy to define by considering a flux through a plaquette $p$ given by $W_{p}=\exp\left(i\frac{2\pi f_{p}}{N}\right)$, where $f_{p}$ is an integer well-defined modulo $N$, for which we choose the unique representative in the range:
$-\frac{N}{2}<f_{p}\leq\frac{N}{2}\;.$ (2)
We have $f_{p}=(ds)_{p}+Nm_{p}$, where $m_{p}$ is some integer defined on plaquettes (a 2-cochain). Further, since $d^{2}=0$ (see e.g. [20, 21]), we have that $(df)_{c}=\sum_{p\in\partial c}f_{p}\equiv 0\bmod N$, with $c$ denoting a cube of the lattice.
In Fig. 1c), we give an example of the center vortex configuration by shading the plaquettes for which $f_{p}\equiv 1\bmod N$. If $(df)_{c}=0$ were to hold, then the center vortex would form closed string worldsheet configurations on the dual lattice. However because $(df)_{c}\equiv 0\mod N$, there can be configurations for which $(df)_{c}=N(dm)_{c}$, where $m_{p}\in\mathbb{Z}$, so $(dm)_{c}\neq 0$ (a loop on the dual lattice) is our definition of a monopole-junction — a worldline where $N$ center-vortex worldsheets meet.
Consider now the $\mathbb{Z}_{7}$ LGT with the usual Wilson action (1). This theory was simulated on the 4-torus that we interpret as the Euclidean space-time of size $L^{4}$. To probe this model, we ran Monte Carlo simulations for 200 values of $\beta\in[0.5,2.7]$, and performed $2\cdot 10^{4}$ measurements each, with 20 decorrelation sweeps between measurements. Furthermore, we chose $L=10$ and estimated the error via bootstrap and present the results in Fig. 2. Note that the errors are too small to be seen.
This theory has three phases, which we can distinguish by the Polyakov loop (i.e. the Wilson loop winding along one of the cycles of the torus). Fig. 2 shows the Polyakov loop expectation value, which shows a clear jump at two values of $\beta_{1}\approx 1.157\pm 0.006$ and $\beta_{2}\approx 2.108\pm 0.006$. The confined phase is clear from the plot, as the Polyakov loop expectation value is (nearly) zero, but the middle (the photon) and the large $\beta$ (the SSB) phases both have a finite Polyakov loop VEV. To distinguish the two, we measure the following. Let $x=(t,x_{1},x_{2},x_{3})$ parametrize the 4D position on a torus (i.e. $t\sim t+L$ and $x_{i}\sim x_{i}+L$). Let $\mathcal{L}_{\bm{x}}$ be the set of all links at the spatial point $\bm{x}=(x_{1},x_{2},x_{3})$ pointing in the temporal direction. Now, define the temporal Polyakov loop to be $P_{\bm{x}}=\prod_{\ell\in\mathcal{L}_{\bm{x}}}U_{\ell}$. For a given configuration, we define the spatially smeared Polyakov loop $\bar{P}=\frac{1}{L^{3}}\sum_{\bm{x}}P_{\bm{x}}$. The $\bar{P}$ is a complex number that depends on the configuration.
In the inset of Fig. 2, we plot the Monte Carlo histogram of this observable. One can clearly see that for the SSB phase, the smeared Wilson loop shows a $\mathbb{Z}_{7}$ symmetric distribution reflecting a spontaneously broken $\mathbb{Z}_{7}$ 1-form symmetry, while in the photon phase, a circular distribution emerges, indicating the emergence of the $\text{U}(1)$ 1-form symmetry. In the confined phase, the distribution of $\bar{P}$ is unsurprisingly sharply peaked around zero. In the inset of Fig. 2, we also show 3D slices of typical configurations in the three phases. The string-like objects are illustrations of the center vortices. One can clearly see that they form short loops (or rather, short worldsheets in 4D) in the $\beta>\beta_{2}$ SSB phase, while they clearly proliferate in both the photon and the confined phase. On the other hand, the monopole-junctions, shown as little cubes on 3D slices, are nonexistent in the SSB phase, are very rare (i.e. form short disconnected loops) in the photon phase, and are abundant in the confined phase. This establishes the connection between center vortices and monopole-junctions as proposed in [7].
This behavior is expected for any $N\geq 5$. However, in [22], a $\mathbb{Z}_{4}$ theory was considered, and there the authors considered a model with an additional parameter, $\tilde{\beta}$ which coupled two charge 2 Wilson loops, i.e. the action was modified to
[…]
They argued that for a particular choice of $\beta$ and $\tilde{\beta}$, a massless photon phase emerges, but their numerics were inconclusive. We computed the smeared Wilson loops and present the results in Fig. 4. We computed the smeared Wilson loops in a 3x8 grid of values $\tilde{\beta}\times\beta\in[-0.379,0.407]\times[1.4,1.6]$ by running $10^{4}$ Monte Carlo sweeps of an $L=6$ lattice. We present the time series histogram of the smeared Wilson loops in Fig. 4. As can be observed, we see the same pattern of circle formation indicating a photon phase in between the confined and the SSB phase, confirming the finding of [22].
Figure 4: The smeared Wilson loop as a function of $\beta$ and $\tilde{\beta}$ (see (3)) of the $\mathbb{Z}_{4}$ lattice gauge theory.
Lastly, we take a look at the $\mathbb{Z}_{3}$ model. In this model with the action (1), there are only two phases. The picture of this model is that as $\beta$ decreases from $\beta\gg 1$ to $\beta\ll 1$, both center vortices and monopole-junctions become abundant simultaneously. To open a photon phase, we therefore suppress the monopoles and write the action
$S=-\beta\sum_{p}\text{Re\;}W_{p}-\mu\sum_{c}(dm)_{c}^{2}\;.$ (4)
To identify a third phase, we look at the action susceptibility (i.e. heat capacity) defined as
$\chi_{S}=\frac{1}{L^{4}}\left(\left\langle S^{2}\right\rangle-\left\langle S\right\rangle^{2}\right).$ (5)
In Fig. 3 we plot numerical Monte Carlo results $\chi_{S}$ as a function of $\beta$ at $\mu=1$. We ran Monte Carlo simulations for $L=8,10,12,$ and $14$ for various values of $\beta$, sampling more densely around the two suspected critical values. Moreover, a monopole suppression of $\mu=1$ was chosen. For each experiment, $4\cdot 10^{4}$ measurements were performed, with 100 decorrelation sweeps between them. The error was estimated using Bootstrap. The results for $\chi_{S}$ can be seen in Fig. 3. As is clear, the susceptibility shows two peaks, indicating a three-phase scenario, with phase transitions at $\beta=\beta_{1}\approx 0.265$ and $\beta=\beta_{2}\approx 0.512$. The inset shows the smeared Polyakov loop $\bar{P}$ for values of beta close to the second transition, showing the circles forming in the photon phase.
Finally, we also consider the connected plaquette-plaquette correlator (PPC) defined as
$C_{x-y;\mu\nu}=\left\langle W_{p_{x,\mu\nu}}^{*}W_{p_{y,\mu\nu}}\right\rangle-\big{\langle}W_{p_{x,\mu\nu}}^{*}\big{\rangle}\big{\langle}W_{p_{y,\mu\nu}}\big{\rangle}\;,$ (6)
where $p_{x,\mu\nu}$ indicates a plaquette with a corner at lattice site $x$, expanding in directions $\mu,\nu=0,1,2,3$, and averages are over ensemble and all spatial translations. At long distances in the photon phase, the theory flows to a pure $\text{U}(1)$ gauge theory with the action $S=\frac{1}{4}\int d^{4}x\;F_{\mu\nu}F^{\mu\nu}$, which describes free photons. All gauge invariant local operators of the $\text{U}(1)$ gauge theory must be built from the field-strength operator $F_{\mu\nu}$, so we expect $C_{x-y;\mu\nu}\propto\left\langle F_{\mu\nu}(x)F_{\mu\nu}(y)\right\rangle$ for large $|x-y|$. Further, we have that
$\left\langle F_{\mu\nu}(x)F_{\rho\sigma}(y)\right\rangle\propto\frac{D_{\mu\nu;\rho\sigma}}{|x-y|^{4}}\;,$ (7)
where $D_{\mu\nu;\rho\sigma}$ is independent of the modulus $|x-y|$ and is easily computed [^4] One way is to note that $A_{\mu}(x)A_{\nu}(y)=\delta_{\mu\nu}/|x-y|^{2}$ in the Feynman gauge, and then take derivatives., but is not relevant for our discussion. Note that if points $x$ and $y$ differ by translation along the time direction $x^{0}$, the correlator $\left\langle F_{ij}(x)F_{ij}(y)\right\rangle>0$, while $\left\langle F_{0i}(x)F_{0i}(y)\right\rangle<0$ by reflection positivity, so it makes sense to look at the positive definite combination of the two, which on the lattice corresponds to
$C(n)=\sum_{\rho}\Big{(}\sum_{\begin{subarray}{c}\mu>\nu\\ \mu,\nu\neq\rho\end{subarray}}C_{n\hat{\rho},\mu\nu}-\sum_{\mu\neq\rho}C_{n\hat{\rho},\rho\mu}\Big{)}\approx\frac{K}{n^{4}}\;,$ (8)
valid up to corrections $\mathcal{O}(1/n^{6})$, where $\hat{\rho}$ is a lattice unit vector in the direction $\rho$, and $K$ is a constant. We ran Monte Carlo simulations of $\mathbb{Z}_{3}$ LGT with $\mu=1$, $\beta=0.5114$ and $L=16$, performing $1.28\cdot 10^{6}$ measurements, with 500 decorrelation sweeps between them. The error was estimated using a binning Jackknife analysis. The quotient $\frac{C(n)}{C(n+1)}$, behaves as $\frac{(n+1)^{4}}{n^{4}}+o(1/n^{2})$ for large $n$ and has no fitting parameter. We compare this theoretical curve with lattice results in Fig. 5 and find excellent agreement, confirming the photon phase of the model.
Figure 5: Lattice data shows the ratio of correlators $C(n)/C(n+1)$ and the smooth red curve shows the theoretical prediction assuming 4th power law decay, consistent with the occurrence of photon phase. Note that the curve has no fitting parameters.
IV Conclusions
We numerically showed that $\mathbb{Z}_{N}$ lattice gauge theories have three phases for all $N\geq 3$, supplementing the spontaneously broken and restored $\mathbb{Z}_{N}$ 1-form symmetry with a phase of emergent photons, where the symmetry enhances to $U(1)$.
In addition, we demonstrated a direct link between the photon phase and the proliferation of string-like center vortices, while the confined phase was linked to proliferation of both vortices and vortex junctions (monopoles) as proposed in [7]. This suggests that center vortices are not directly responsible for confinement as commonly believed for $SU(N)$ gauge theory. [^5] Defining monopoles and center vortices requires gauge fixing in $SU(N)$ gauge theories (see e.g. [23, 10, 11]), although there are some efforts to identify such objects without it [24, 24]. Further, a novel proposal of how to discretize $\text{SU}(N)$ gauge theories [25, 26], may be related to a gauge invariant definition of vortices and monopoles [27].
Our work indicates that the Landau paradigm for $\mathbb{Z}_{N}$ 1-form symmetries must include a massless phase when $N\geq 3$, distinguishing it from its 0-form counterpart.
The case of $\mathbb{Z}_{2}$ lattice gauge theory is more subtle as it does not have a global charge conjugation symmetry $C$ [^6] TS would like to thank M. Nguyen and E. Richards for discussions on this point. While it was proposed in [7] to have a photon phase when deformed, such deformations would introduce the $C$-symmetry. A similar issue occurs in $D_{N}$ lattice gauge theories, which have no $C$ symmetry and only at most a $\mathbb{Z}_{2}$ 1-form symmetry. Yet for large enough $N$, $D_{N}$ LGT should be indistinguishable from $O(2)$ LGT, which certainly has a photon phase, albeit with a non-invertible emergent 1-form symmetry. The Landau paradigm for $\mathbb{Z}_{2}$ symmetry therefore likely involves a massless phase as suggested in [7], although the details are more subtle and are left for future explorations.
Acknowledgement
We would like to thank Mathew Bullimore, Xavier Crean, Christof Gattringer, Tyler Helmuth, Nabil Iqbal, Biagio Lucini, Mendel Nguyen, Emily Richards, Shu-Heng Shao and Yuya Tanizaki for the discussions and comments. This work has made use of the Hamilton HPC Service of Durham University. TS is supported by the Royal Society University Research Fellowship and, in part, by the STFC consolidated grant number ST/T000708/1. JG is supported by EPSRC grant EP/Y028872/1.